OpenStudy (anonymous):

Words don't belong in math >.< Amy has a job transporting soft drinks by truck. Her truck is filled with cans that weigh 14 ounces and bottles that weigh 70 ounces. There is a combined total of 930 cans and bottles in her truck. Letting x be the number of the 14-ounce cans in her truck, write an expression for the combined total weight (in ounces) of the cans and bottles in her truck.

OpenStudy (amistre64):

930 = 14x + 70y maybe?

OpenStudy (saifoo.khan):

Lets read.

OpenStudy (amistre64):

total number, not total weight

OpenStudy (amistre64):

x+y = 930 maybe :)

OpenStudy (anonymous):

heh I have no idea >.<

OpenStudy (amistre64):

i think ive confused me self lol

OpenStudy (anonymous):

I was thinkg 14x+70y=930 but like you said.. cans.. not weight... grrr

OpenStudy (amistre64):

Letting x be 14-ounce cans in her truck, write an expression for the combined total weight (in ounces) of the cans and bottles in her truck 14x + 70y = 930(z)

OpenStudy (amistre64):

that dont make sense either ...

OpenStudy (anonymous):

haha see why I hate words in math??

OpenStudy (amistre64):

x + y = 930 y = 930 -x 14(x) + 70 (930-x) = total weight

OpenStudy (amistre64):

simplificate as wanted :)

OpenStudy (anonymous):

heh look good Thank you kindly :D

OpenStudy (anonymous):

Yep. It was a system of equations question. This is one of those things that looks impossible on a computer, but when written on paper is a piece of pi.